OpenStudy (anonymous):

Solve. 3x^2 – 5x + 7 = 0 x= -5 +_ sqrt 109 / 6 x=5 +_ sqrt 109 / 6 x= -5 +_ i sqrt 59 / 6 x= 5 +_ i sqrt 59 /6

OpenStudy (anonymous):

x= (5 +_ i sqrt 59 )/6\[\left\{x\to \frac{1}{6} \left(5-i \sqrt{59}\right),x\to \frac{1}{6} \left(5+i \sqrt{59}\right)\right\} \]